Please solve, show work, and give detail explanation4. A frim has bonds with 13 years to maturity, a yield-to-maturity (YTM) of 9.2 percent, and a current price of $928.81. The
Answer #1

1.
Using financial calculator
N=13*2=26
I/Y=9.2%/2=4.6%
PV=-928.81
FV=1000
CPT PMT=41.25

Hence, coupon rate=41.25*2/1000=8.250%


2.
=Clean price+days since last coupon/period between 2 coupons*periodic coupon
=945+4/6*7.3%*1000/2
=969.33